one angle in a triangle is half the largest angle but three times the smallest. find all three angles.
Answer by mananth(16946)

let the angle be x
largest is 2x
smallest = x/3
..
x+2x+x/3 =180
3x+6x+x / 3 =180
10x=540
x=54 one angle
largest = 2x= 2*54=108
smallest = x/3 = 18
...
CHECK
54+18+108=180